The demand function Q(P) and cost functions C(Q) of a company's are given by the equations:

Q=12000−60P

(Q)=10000+4Q,

where P and Q are the price and quantity, respectively.

What is the company's profit function?




a.Profit=−60P−4Q+2 000


b.Profit=−60P^2+11 760P−58 000


c.Profit=−60P^2+12 240P−58 000


d.Profit=−60P^2+12 240P+38 000


1
Expert's answer
2022-04-25T17:24:08-0400

 What is the company's profit function in terms P?

(Q) = 10 000 + 4Q = 10000 + 4(12000 - 60 P) = 58000 - 240 P

Profit = TR - Q= 12 000 P - 60 P2 - (58 000 - 240 P) = -60 P2 + 12240P - 58 000

Aswer: c